What are the zero(s) for this equation:
4(x+6)^2 - 8 = -4
a. x = {-5}
b. x = {-5,-7}
c. x = {-5,0}
d. x = {5,-7}

Answer by nabla(475)

Divide both sides by 4, giving:
(x+6)^2-2=-1
(x+6)^2=1
x+6=+/- 1
x=-6+/-1
-6+1=-5
-6-1=-7
Giving (b) as the answer.